OpenStudy (anonymous):

(4 to the 3rd) to the 5th

OpenStudy (celinegirl):

Hi! welcome to Open Study! What is your question?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle (4^3)^5 }\). That's the question.

OpenStudy (solomonzelman):

Well, that would be a very huge number if you want to calculate it numerically, however, it is also possible to write it in terms of base 4. \(\color{#000000 }{ \displaystyle (4^3)^5=4^{3\times 5}=4^{~?~} }\) (you tell me the exponent)